OpenStudy (anonymous):

The positive difference between two consecutive even perfect squares is 268 . Compute the larger of the two squares.

OpenStudy (anonymous):

so it is (x+1)^2-x^2=268

OpenStudy (anonymous):

says even right, so maybe \[(2n+2)^2-(2n)^2\]

OpenStudy (anonymous):

x^2+1^2-x^2=268 x^2+1-x^2=268 x^2-x^2=268-1 0=267 I did something wrong....

OpenStudy (anonymous):

oh, I see

OpenStudy (anonymous):

2n^2+4-2n^2=268 2n^2-2n^2=264

OpenStudy (anonymous):

0=264?!?!

OpenStudy (anonymous):

i made a mistake

OpenStudy (anonymous):

\[(2n+2)^2-(2n)^2=8n+4\]

OpenStudy (anonymous):

i forgot to square the 2!!

OpenStudy (anonymous):

where didya get the "8n"?

OpenStudy (anonymous):

\[8n+4=268\] \[n=33\] \[2\times 33+2\] is your answer

OpenStudy (anonymous):

algebra

OpenStudy (anonymous):

\[(2n+2)(2n+2) -4n^2\] etc